OpenStudy (anonymous):

:) Use the quadratic formula to solve the problem. A soccer player throws the ball onto the field from the sideline. An equation modeling the ball's flight is h = -16t2 + 18t + 2. In how many seconds will the ball reach the ground? Choose one answer. a. in about 2 seconds b. in about 1.2 seconds c. in about 1.8 seconds d. Cannot be determined

OpenStudy (johnweldon1993):

So using your quadratic formula with -16 for 'a' 18 for 'b' and 2 for 'c' plug those into \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] you will get 2 times...1 of which will most likely be negative....but you can ignore that...since we only want the positive number

OpenStudy (johnweldon1993):

edit * you will get 2 results!*

OpenStudy (reemii):

 yes, the ball touches the ground when h(t)=0.

OpenStudy (anonymous):

-4/32 or -1/8 ?

OpenStudy (anonymous):

wait thats not even an option

OpenStudy (johnweldon1993):

not quite...try that again the 2 results I get are 1.22 and -.101

OpenStudy (reemii):

john, well done! xp

OpenStudy (anonymous):

So how did i not get that. Wtf ugg oh well. Thanks

OpenStudy (johnweldon1993):

well no try it again though \[\frac{ -18\pm \sqrt{18^2 - 4(-16)(2)} }{ 2(-16) }\] is what you should have done

OpenStudy (anonymous):

that is what i did.

OpenStudy (reemii):

there's a \(\pm\) in the formula. it is the "-" that is giving the answer 1.2

OpenStudy (johnweldon1993):

so then lets simplify whats under that radical 18² is 324 -4 times -16 times 2 is 128 so now you have \[\frac{ -18\pm \sqrt{452} }{ -32 }\]

OpenStudy (johnweldon1993):

*It's √452 because you add 324 and 128*

OpenStudy (johnweldon1993):

so now you evaluate the √452 ...which will give you 21.2602916254 so you have 2 equations \[\frac{ -18 + 21.2602916254 }{ -32 }\] and \[\frac{ -18 - 21.2602916254 }{ -32 }\]

OpenStudy (johnweldon1993):

so do those out and what are the 2 results you get?

OpenStudy (johnweldon1993):

Well...she left...regardless....There's the work :D lol